# Client Selection for Federated Policy Optimization with Environment Heterogeneity

### v6

@@ Appendix G Proof of Lemma 12
$\mathbb{E}_{\text{init}}\left|\left[u_{\vartheta}(s,a)-u_{\vartheta^{\prime}}(s,a^{\prime})\right|\right]^{2}$ $\leq 2\mathbb{E}_{\text{init}}\left[\left(\sum_{i}^{m}\left([\vartheta]_{i}^{T}(s,a)-[\vartheta^{\prime}]_{i}^{T}(s,a^{\prime})\right)^{2}\right)+\left\|\vartheta^{\prime}\right\|_{2}^{2}\right]$
$\leq 2\mathbb{E}_{\text{init}}\left[\left(\sum_{i}^{m}2\left([\vartheta]_{i}^{T}(s,a)\right)^{2}+2\left([\vartheta^{\prime}]_{i}^{T}(s,a^{\prime})\right)^{2}\right)+\left\|\vartheta^{\prime}\right\|_{2}^{2}\right]$
@@
Proof  By the triangle inequality and Lemma 18, we have
$\bar{\varepsilon}_{w}$ $=\max_{t}\left\|V^{t}-\bar{V}^{t}\right\|$
$=\max_{t}\mathbb{E}_{\text{init}}\left[\left\|V^{t}-\bar{V}^{t}\right\|\right]$
$\leq\max_{t}\mathbb{E}_{\text{init}}\left[\left|V^{t}(s)-V^{t,0}(s)\right|+\left|V^{t,0}(s)-\bar{V}^{t}(s)\right|\right]$
$=\max_{t}\mathbb{E}_{\text{init}}\left[\left|V^{t}(s)-V^{t,0}(s)\right|+\left|\sum_{n=1}^{N}q_{n}V_{n}^{t,0}(s)-\sum_{n=1}^{N}q_{n}V_{n}^{t}(s)\right|\right]$
@@
Let $\left(t^{\prime},s^{\prime}\right)=\arg\max_{t>0,s\in\mathcal{S}}\left\|\pi^{t}(\cdot|s)-\sum_{n=1}^{N}q_{n}\pi^{t}_{n}(\cdot|s)\right\|_{2}$, we have
@@
$\quad\left.\left.+\left|\log\frac{\sum_{a^{\prime}\in\mathcal{A}}\exp\left(f_{\theta^{t}}(s^{\prime},a^{\prime})\right)}{\sum_{a^{\prime}\in\mathcal{A}}\exp\left(f_{\theta^{t}_{n}}(s^{\prime},a^{\prime})\right)}\right|\right)\right)^{1/2}.$ (70)
@@
$\leq\left(2\max_{a\in\mathcal{A},n\in[N]}\left(\left|f_{\theta^{t^{\prime}}}(s^{\prime},a)-f_{\theta^{t^{\prime}}_{n}}(s^{\prime},a)\right|+\left|\log\frac{\sum_{a^{\prime}\in\mathcal{A}}\exp\left(f_{\theta^{t}}(s^{\prime},a^{\prime})\right)}{\sum_{a^{\prime}\in\mathcal{A}}\exp\left(f_{\theta^{t}_{n}}(s^{\prime},a^{\prime})\right)}\right|\right)\right)^{1/2}.$ (71)
@@
$\max_{a\in\mathcal{A},n\in[N]}\left|f_{\theta^{t^{\prime}}}(s^{\prime},a)-f_{\theta^{t^{\prime}}_{n}}(s^{\prime},a)\right|$ $=\mathcal{O}\left(R_{\theta}\right).$ (72)
@@
$\log\sum_{a^{\prime}\in\mathcal{A}}\exp\left(f_{\theta^{t}}(s,a^{\prime})\right)-\log\sum_{a^{\prime}\in\mathcal{A}}\exp\left(f_{\theta^{t}_{n}}(s,a^{\prime})\right)$ $\leq\max_{a\in\mathcal{A}}f_{\theta^{t}}(s,a)-\frac{1}{\left|\mathcal{A}\right|}\sum_{a^{\prime}}\left(f_{\theta_{n}^{t}}(s,a^{\prime})\right)$
$=\mathcal{O}\left(R_{\theta}\right),$
@@
$\log\sum_{a^{\prime}\in\mathcal{A}}\exp\left(f_{\theta^{t}_{n}}(s,a^{\prime})\right)-\log\sum_{a^{\prime}\in\mathcal{A}}\exp\left(f_{\theta^{t}}(s,a^{\prime})\right)$ $\leq\max_{a\in\mathcal{A}}f_{\theta_{n}^{t}}(s,a)-\frac{1}{\left|\mathcal{A}\right|}\sum_{a^{\prime}}\left(f_{\theta^{t}}(s,a^{\prime})\right)$
$=\mathcal{O}\left(R_{\theta}\right),$
@@
$\max_{a\in\mathcal{A},n\in[N]}\left|\log\frac{\sum_{a^{\prime}\in\mathcal{A}}\exp\left(f_{\theta^{t}}(s,a^{\prime})\right)}{\sum_{a^{\prime}\in\mathcal{A}}\exp\left(f_{\theta^{t}_{n}}(s,a^{\prime})\right)}\right|$ $=\mathcal{O}\left(R_{\theta}\right).$ (73)
@@
$\bar{\varepsilon}_{\theta}$ $=\mathcal{O}\left(R_{\theta}^{1/2}\right),$

### v7

@@ Appendix G Proof of Lemma 12
$\mathbb{E}_{\text{init}}\left|\left[u_{\vartheta}(s,a)-u_{\vartheta^{\prime}}(s,a^{\prime})\right|\right]^{2}$ $\leq 2\mathbb{E}_{\text{init}}\left[\left(\sum_{i}^{m}\left(\vartheta_{i}^{T}(s,a)-\left(\vartheta^{\prime}_{i}\right)^{T}(s,a^{\prime})\right)^{2}\right)+\left\|\vartheta^{\prime}\right\|_{2}^{2}\right]$
$\leq 2\mathbb{E}_{\text{init}}\left[\left(\sum_{i}^{m}2\left(\vartheta_{i}^{T}(s,a)\right)^{2}+2\left(\left(\vartheta^{\prime}_{i}\right)^{T}(s,a^{\prime})\right)^{2}\right)+\left\|\vartheta^{\prime}\right\|_{2}^{2}\right]$
@@
Proof  Let $(t^{\prime},s^{\prime})=\arg\max_{t>0,s\in\mathcal{S}}\left|V^{t}(s)-\bar{V}^{t}(s)\right|$, which are dependent on the initialization $w^{0}$. By the triangle inequality and Lemma 18, we have
$\mathbb{E}_{\text{init}}\left[\bar{\varepsilon}_{w}\right]$ $=\mathbb{E}_{\text{init}}\left[\max_{t}\left\|V^{t}-\bar{V}^{t}\right\|\right]$
$\leq\mathbb{E}_{\text{init}}\left[\left|V^{t^{\prime}}(s^{\prime})-V^{t^{\prime},0}(s^{\prime})\right|+\left|V^{t^{\prime},0}(s^{\prime})-\bar{V}^{t^{\prime}}(s^{\prime})\right|\right]$
$=\mathbb{E}_{\text{init}}\left[\left|V^{t^{\prime}}(s^{\prime})-V^{t^{\prime},0}(s^{\prime})\right|+\left|\sum_{n=1}^{N}q_{n}V_{n}^{t^{\prime},0}(s^{\prime})-\sum_{n=1}^{N}q_{n}V_{n}^{t^{\prime}}(s^{\prime})\right|\right]$
@@
Let $\left(t^{\prime},s^{\prime}\right)=\arg\max_{t>0,s\in\mathcal{S}}\left\|\pi^{t}(\cdot|s)-\sum_{n=1}^{N}q_{n}\pi^{t}_{n}(\cdot|s)\right\|_{2}$, which is conditional on $\theta^{0}$, we have
@@
$\quad\left.\left.+\left|\log\frac{\sum_{a^{\prime}\in\mathcal{A}}\exp\left(f_{\theta^{t}}(s^{\prime},a^{\prime})\right)}{\sum_{a^{\prime}\in\mathcal{A}}\exp\left(f_{\theta^{t}_{n}}(s^{\prime},a^{\prime})\right)}\right|\right)\right)^{1/2}.$
@@
$\leq\left(2\max_{a\in\mathcal{A},n\in[N]}\left(\left|f_{\theta^{t^{\prime}}}(s^{\prime},a)-f_{\theta^{t^{\prime}}_{n}}(s^{\prime},a)\right|+\left|\log\frac{\sum_{a^{\prime}\in\mathcal{A}}\exp\left(f_{\theta^{t}}(s^{\prime},a^{\prime})\right)}{\sum_{a^{\prime}\in\mathcal{A}}\exp\left(f_{\theta^{t}_{n}}(s^{\prime},a^{\prime})\right)}\right|\right)\right)^{1/2}.$
By Jensen’s inequality, we have
$\mathbb{E}_{\text{init}}\left[\bar{\varepsilon}_{\theta}\right]$ $\leq\mathbb{E}_{\text{init}}\left[2\max_{a,n}\left|f_{\theta^{t^{\prime}}}(s^{\prime},a)-f_{\theta^{t^{\prime}}_{n}}(s^{\prime},a)\right|+2\left|\log\frac{\sum_{a^{\prime}\in\mathcal{A}}\exp\left(f_{\theta^{t}}(s^{\prime},a^{\prime})\right)}{\sum_{a^{\prime}\in\mathcal{A}}\exp\left(f_{\theta^{t}_{n}}(s^{\prime},a^{\prime})\right)}\right|\right]^{1/2}.$ (70)
@@
$\mathbb{E}_{\text{init}}\left[\max_{a\in\mathcal{A},n\in[N]}\left|f_{\theta^{t^{\prime}}}(s^{\prime},a)-f_{\theta^{t^{\prime}}_{n}}(s^{\prime},a)\right|\right]$ $=\mathcal{O}\left(R_{\theta}\right).$ (71)
@@
$\log\sum_{a^{\prime}\in\mathcal{A}}\exp\left(f_{\theta^{t}}(s,a^{\prime})\right)-\log\sum_{a^{\prime}\in\mathcal{A}}\exp\left(f_{\theta^{t}_{n}}(s,a^{\prime})\right)$ $\leq\max_{a\in\mathcal{A}}f_{\theta^{t}}(s,a)-\frac{1}{\left|\mathcal{A}\right|}\sum_{a^{\prime}}\left(f_{\theta_{n}^{t}}(s,a^{\prime})\right),$
@@
$\log\sum_{a^{\prime}\in\mathcal{A}}\exp\left(f_{\theta^{t}_{n}}(s,a^{\prime})\right)-\log\sum_{a^{\prime}\in\mathcal{A}}\exp\left(f_{\theta^{t}}(s,a^{\prime})\right)$ $\leq\max_{a\in\mathcal{A}}f_{\theta_{n}^{t}}(s,a)-\frac{1}{\left|\mathcal{A}\right|}\sum_{a^{\prime}}\left(f_{\theta^{t}}(s,a^{\prime})\right),$
@@
$\mathbb{E}_{\text{init}}\left[\max_{a\in\mathcal{A},n\in[N]}\left|\log\frac{\sum_{a^{\prime}\in\mathcal{A}}\exp\left(f_{\theta^{t}}(s,a^{\prime})\right)}{\sum_{a^{\prime}\in\mathcal{A}}\exp\left(f_{\theta^{t}_{n}}(s,a^{\prime})\right)}\right|\right]$ $=\mathcal{O}\left(R_{\theta}\right).$ (72)
@@
$\mathbb{E}_{\text{init}}\left[\bar{\varepsilon}_{\theta}\right]$ $=\mathcal{O}\left(R_{\theta}^{1/2}\right),$